# HeadCraft: Modeling High-Detail Shape Variations for Animated 3DMMs

### v1

@@ B.1 Applications
$\hskip 64.01869ptg(map(z))\cdot(1-M))$
$+\lambda^{\textrm{face}}\|g(map(z)+\Delta w)\cdot M^{\textrm{face}}\|_{1},$
@@
The third and second “preserve” terms are introduced to not let the map guided by the $\Delta w$ optimization deviate much from the output corresponding to the regressed $z$ in non-visible regions, which is essential due to the tendency of convergence to the average shape there when optimizing in the $\mathcal{W}+$ space. $\lambda^{\textrm{LPIPS}}=0.1$ and $\lambda^{L_{1}}=3$ remain the same as before, and $\lambda_{\textrm{preserve}}^{\textrm{LPIPS}}=0.01$ and $\lambda_{\textrm{preserve}}^{L_{1}}=0.3$ are selected 10×10$\times$10 × less. The last regularization term is introduced to avoid hallucinations in the facial region that otherwise become visually apparent in the non-visible region of the map even in case of relatively small high-frequency inconsistency. $\lambda^{\textrm{face}}$ is set to $\frac{10}{256}$. The optimization is carried out by Adam with the same learning rate of $10^{-2}$ for 500 steps. The $\Delta w$ is initialized with zeros.
@@
Figure 13: Additional results of the diversity and level of detail of the unconditionally sampled generations from HeadCraft. The generations are obtained by randomly sampling $z\sim\mathcal{N}(0,\mathbb{I})$ latent code of the generative model. The displacements, returned by the model, are applied to the random FLAMEs sampled from Gaussian distribution with statistics calculated over the NPHM dataset.

### v2

@@ B.1 Applications
Figure 14: An example of the use case for the HeadCraft registration and fitting stages to Kinect depth data. An example of a frontal RGB frame is shown on the left (the color information is not used further), along with the observed Kinect depth, converted into a mesh and aligned with the standard FLAME coordinate system. We fit FLAME and apply the partial registration procedure, described in Sec. B, to register the displacements in the observed region (middle) and further fit the HeadCraft latent representation to the corresponding part of the UV map.
@@
$\hskip 64.01869ptg(map(z))\cdot(1-M)),$
@@
The third and second “preserve” terms are introduced to not let the map guided by the $\Delta w$ optimization deviate much from the output corresponding to the regressed $z$ in non-visible regions, which is essential due to the tendency of convergence to the average shape there when optimizing in the $\mathcal{W}+$ space. $\lambda^{\textrm{LPIPS}}=0.1$ and $\lambda^{L_{1}}=3$ remain the same as before, and $\lambda_{\textrm{preserve}}^{\textrm{LPIPS}}=0.01$ and $\lambda_{\textrm{preserve}}^{L_{1}}=0.3$ are selected 10×10$\times$10 × less. The optimization is carried out by Adam with the same learning rate of $10^{-2}$ for 500 steps. The $\Delta w$ is initialized with zeros.
@@
Use case: Kinect data. We demonstrate that the HeadCraft pipeline can be applied to real-world depth scans in Fig. 14. For a sample RGB-D image captured by Kinect, we first convert it into a mesh by unprojecting the points with color and connecting the vertices by the triangles constructed from the image pixels. The resulting mesh with vertex colors is rendered onto three views (frontal, slight left, and slight right) to obtain the facial landmarks from each side via an image-based facial landmark detector [8] and aggregate them (jawline landmarks are obtained from slight left and slight right and the others from the frontal rendering). FLAME is fitted using these landmarks, and the displacements for the visible part of the mesh are obtained via the partial registration procedure described above and later baked into the UV displacement map. We also show the result of fitting the latent representation of HeadCraft to the visible part of the scan. Compared to the aforementioned procedure, we omit LPIPS loss to give L1 more relevance in predicting coarse shape and only supervise the latent in the scalp region.
@@
Figure 15: Additional demonstration of the diversity and level of detail of the unconditionally sampled generations from HeadCraft. The generations are obtained by randomly sampling $z\sim\mathcal{N}(0,\mathbb{I})$ latent code of the generative model. The displacements, returned by the model, are applied to the random FLAMEs sampled from Gaussian distribution with statistics calculated over the NPHM dataset.